# Lrq-Fact: LLM-Generated Relevant Questions for Multimodal Fact-Checking

### v1

@@ 3.4 Rule-Based Decision-Maker
In the final stage of our framework, the decision-maker integrates insights from both the textual and visual analyses to make a conclusive determination about the authenticity of the content. This stage is crucial for verifying the consistency between the image and text, and for identifying any misinformation. The decision-making process is guided by a rule-based module, designed to mimic the judgment process similar to the human fact-checkers. The LLM combines the information gathered from image and text analysis to evaluate potential inconsistencies or factual inaccuracies. The LLM will compare the image description generated by the VLM with the claims made in the news article. For example, if the article claims that a peaceful protest occurred but the image description shows signs of violence, this discrepancy is flagged. Similarly, if the article mentions a political figure being present but the image does not include this individual, the decision-maker identifies this as a significant inconsistency. This cross-modal evaluation ensures that the image and text together do not create a misleading or deceptive narrative.
The rule-based module follows a systematic process:
  • Step 1: It checks whether any of the factual claims in the text can be objectively refuted. If there is credible evidence contradicting a claim, the decision-maker labels it as “Textual Veracity Distortion”.
  • Step 2: If the text is not refuted, it evaluates the image for discrepancies. If the image contains elements that contradict facts or appear manipulated, the label “Visual Veracity Distortion” is applied.
  • Step 3: If both the text and image are individually accurate, it assesses whether they are consistent with one another. A mismatch between the text and image, such as a beach being described while the image shows a city, results in a “Mismatch” label.
  • Step 4: If none of the above issues are found, the content is labeled as “Real”.
Moreover, by providing additional guidelines ensure that the process is thorough:
  • Overall Consistency: It evaluates whether the main ideas, actions, and elements in the article align with the image description. Minor discrepancies are acceptable unless they change the overall meaning.
  • Visual Cues: Details such as body language, facial expressions, and environmental features are examined to ensure alignment with the article’s description.
  • Unrealistic Elements: It checks for AI-generated or unnatural elements in the image. Such details may indicate the image is fabricated, although the decision-maker weighs these findings based on their relevance to the article’s core narrative.
  • Final Judgment: It makes a judgment based on the overall consistency, details, and realism of both the image and the article. The decision is based on whether the story is coherent and plausible, or if contradictions and discrepancies raise doubts. Importantly, it does not over-rely on suggestions of manipulation from generated answers but makes a balanced decision based on the bigger picture.
  • Select the Most Relevant Label: After analysis, it assigns one of the following labels:
    • Textual Veracity Distortion: The text contains false or misleading claims.
    • Visual Veracity Distortion: The image description includes fabricated or unrealistic elements.
    • Mismatch: The text and image depict unrelated events or contexts.
    • Real: The text and image are consistent, without contradictions or unrealistic details.
  • Provide a Detailed Explanation: It outlines key factors influencing its judgment, focusing on text-image consistency and referencing relevant questions and answers. If generated answers hint at manipulation, the decision-maker acknowledges this but ensures it does not dominate the decision if other factors play a more significant role.
Besides providing the final judgment, the decision-maker module also provides a clear and detailed rationale explaining why the content was classified as real or false. For instance, if misinformation is detected, the decision-maker will clarify whether the decision was influenced by issues in the visual content, inaccuracies in the text, or inconsistencies between the two, based on the generated QAs. This transparent reasoning enhances trustworthiness in the fact-checking process by offering insights into how the conclusions were reached.
@@
For each piece of analyzed content, the decision-maker provides: ❶ a final judgment $j_{i}$, such as “Real” or “Fake”, ❷ a label $l_{i}$ identifying the specific issue: “Textual Veracity Distortion”, “Visual Veracity Distortion”, or “Mismatch” and ❸ a detailed explanation of the decision-making process $e_{i}$. This rule-based approach ensures that the framework provides clear, evidence-based conclusions, carefully weighing the alignment between textual and visual information to detect misinformation. The abovementioned process can be formulated as:
(3.6) $qa_{i}^{v}=\oplus_{j=1}^{m}[ques_{i}^{v_{j}},ans_{i}^{v_{j}}],$
(3.7) $qa_{i}^{t}=\oplus_{j=1}^{n}[ques_{i}^{t_{j}},ans_{i}^{t_{j}}],$
(3.8) $j_{i},l_{i},e_{i}={\rm LLM}(text_{i},img_{i},des_{i},qa_{i}^{v},qa_{i}^{t}),$
where $\oplus$ represents the concatenation operation.
4 Experiments
In this section, we conduct extensive experiments on two benchmarks to show that our framework improves detection accuracy and rationale quality for multimodal misinformation compared to existing approaches. Additionally, we assess the framework’s generalizability across various model backbones and evaluate the robustness of different prompting strategies, providing key insights for further optimization and refinement.
@@ 4.1 Datasets
We used two multimodal misinformation detection benchmarks to evaluate the performance of our framework.
MMFakeBench: MMFakeBench dataset [45] contains 11,000 image-text pairs, specifically designed to evaluate misinformation detection models across a broad range of scenarios. MMFakeBench goes beyond the assumption of single-source forgery and presents samples with “Real”, “Textual Veracity Distortion”, “Visual Veracity Distortion”, and “Cross-modal Consistency Distortion”, making it a more realistic and comprehensive benchmark for assessing multimodal misinformation detection systems. This benchmark is particularly useful in testing the capabilities of models like ours that aim to detect inconsistencies between text and image modalities.
DGM4: DGM4 dataset [61] is a recently released large-scale multimodal manipulation dataset comprising 230,000 image-text paired samples. DGM4 includes over 77,000 pristine pairs and 152,000 manipulated pairs, with manipulations in both image and text modalities. Image manipulation in the dataset involves “face swapping and facial emotion editing”, while text manipulation includes “sentence replacement and textual sentiment editing”. The DGM4 dataset is constructed based on the VisualNews dataset [41], which collects data from multiple news agencies like BBC, The Guardian, USA TODAY, and The Washington Post. This diversity introduces significant distribution discrepancies, making DGM4 ideal for testing the model’s performance under open-world, domain-shift scenarios. The dataset is partitioned into four subsets based on news sources, allowing the evaluation of model robustness in handling different writing styles and regional perspectives.
Table 1: Performance comparison (%) of different models for multi-class classification on the MMFakeBench and DGM4 datasets. The last two highlighted rows represent the results of Lrq-Fact. Bold: best result; Underline: second best result.
Vision-Language Language MMFakeBench (N=1000) DGM4 (N=1000)
@@
Language Model with 7B Parameter
InstructBLIP [11] Vicuna-7B 7.1 7.9 6.5 7.8 5.3 11.5 4.6 4.5
VILA [39] Llama2-7B 11.5 7.5 25.0 30.0 18.2 27.2 19.0 19.8
BLIP-2 [34] FlanT5-XL 16.4 20.1 27.5 33.0 20.6 32.9 20.7 21.8
LLaVA-1.6 [42] Vicuna-7B 17.4 14.8 25.7 30.8 19.6 30.5 19.0 19.8
Language Model with 13B Parameter
VILA [39] Llama2-13B 11.5 7.5 25.0 30.0 19.4 27.5 19.0 19.8
InstructBLIP [11] Vicuna-13B 13.7 13.2 24.0 28.8 19.2 27.9 19.0 19.8
BLIP-2 [34] FlanT5-XXL 16.7 34.9 27.3 32.8 18.1 26.3 17.9 19.1
LLaVA-1.6 [42] Vicuna-13B 12.0 22.5 25.0 30.0 19.9 29.7 20.1 22.6
Language Model with 34B Parameter
LLaVA-1.6 [42] Nous-Hermes-2-Yi-34B 25.7 44.5 33.7 40.4 32.5 34.3 36.3 39.4
ChatGPT and Llama3.1
GPT-4V-1.7T [53] ChatGPT 51.0 66.8 49.7 54.0 50.6 51.6 55.9 58.5
Paligemma-3B [6] Llama3.1-70B 61.5 67.7 59.1 61.0 54.4 55.1 60.1 60.9
Paligemma-3B [6] ChatGPT 63.3 65.1 61.2 64.3 56.4 57.4 62.1 62.5
@@ 4.2 Experiment Details and Settings
For MMFakeBench, we use a sample of 1,000 instances from the validation set, with 300 labeled as real, 300 as textual veracity distortion, 100 as visual veracity distortion, and 300 as cross-modal consistency distortion. A similar sampling approach is applied to the DGM4 dataset, ensuring the same number of samples for each manipulation type. This balanced approach across the different categories ensures a fair comparison between domains.
@@
Our framework integrates both LLMs and VLMs to handle the tasks of question generation and answering. For generating image-focused and text-focused questions, we use the Llama 3.1 [71] model with 70 billion parameters. For each modality, we generate five questions. Rather than fine-tuning Llama 3.1 for this task, we employ in-context learning, where 10 example questions are provided for each modality to guide the model in generating precise, high-quality and relevant questions. To answer the image-focused questions, we leverage Paligemma [6], Google’s vision-language model with 3 billion parameters, which excels at object detection and generating detailed image descriptions. Paligemma’s ability to capture intricate visual contexts makes it particularly effective for visual content analysis. To answer text-focused questions, we employ GPT-4o [3], which is equipped with an extensive knowledge base. This allows the model to provide reliable, well-informed answers, ensuring the accurate verification of textual content. In the decision-making phase, GPT-4o is also used to evaluate the generated question-answer pairs. Known for its strong reasoning and rule-following capabilities, GPT-4o assesses the veracity of the content and offers clear rationales for its conclusions. To further explore the versatility of our framework, we also experiment with using Llama 3.1 70B as both the question-answering model and the decision-maker. All experiments are conducted on NVIDIA 40GB V100 GPUs, providing the computational power required for these extensive evaluations.
4.3 Evaluation Metrics
@@
To assess the performance of the various baselines, we employ a multi-class classification approach. Following the practices established in prior works [86, 56, 9], we use the widely adopted macro-F1 score as our primary evaluation metric. The macro-F1 score provides a balanced measure of precision and recall through their harmonic mean, ensuring fair evaluation across all classes. In addition to the macro-F1 score, we also report macro-precision, macro-recall, and macro-accuracy as complementary metrics, offering a more comprehensive understanding of model performance.
4.4 Comparison Models
@@
We select a diverse range of VLMs with different model sizes as baseline models for comparison. These include: (i) 7B and 13B-parameter models, such as InstructBLIP [11], VILA [39], BLIP-2 [34], and LLaVA-1.6 [42]; (ii) a 34B-parameter model, also represented by LLaVA-1.6; and (iii) the closed-source GPT-4V-1.7T [53] model.
4.5 Experiment Results
We conducted extensive experiments to assess the performance of our proposed framework, Lrq-Fact, in comparison to baseline models on the MMFakeBench and DGM4 datasets.
4.5.1 Main Results.
We provide a comprehensive comparison of various baseline models using the MMFakeBench and DGM4 datasets, as shown in Table 1. Our Lrq-Fact framework employs two different LLMs as decision-makers: GPT-4o, which achieves an F1-score of 63.3%, and Llama-3.1-70B, with an F1-score of 61.5% on MMFakeBench—both outperforming GPT-4V. Notebly, despite having fewer parameters than GPT-4V, our framework consistently delivers better performance, underscoring its efficiency and effectiveness in handling multimodal misinformation detection. Among open-source models, LLaVA-1.6-34B is the leading performer, yet it only achieves an F1-score of 25.7%, significantly trailing GPT-4V, which scores 51%. This reveals a substantial gap in detection capabilities between open-source and proprietary models. Within the same model family, such as LLaVA-1.6-Vicuna-7B and LLaVA-1.6-34B, models with more parameters consistently demonstrate better performance. Smaller VLMs struggle with instruction-following and maintaining high predicted consistency, suggesting that models with 7B parameters lack the multimodal comprehension required to effectively combat misinformation. Due to these limitations in smaller models, we selected mid-sized open-source and proprietary models as baselines for comparison against Lrq-Fact. Further evaluation results for binary classification are provided in the Appendix  A.1
4.5.2 Comparison Against Single-Source Detectors.
We compare Lrq-Fact framework with LLaVA-1.6-34B, GPT-4V-1.7T, and several state-of-the-art single-source detectors [21, 44, 61]. These detectors are representative of the top-performing methods for single-source misinformation detection. Details on each detector are provided in the Appendix  A.2. To ensure a fair comparison, beyond evaluating existing single-source detectors, we also integrate three of the most powerful detectors—FakeingFakeNews [21], LNP [40], and HAMMER [61]—to assess the performance of Lrq-Fact. The results, summarized in Table 2, demonstrate that Lrq-Fact outperforms LLaVA-1.6-34B, GPT-4V, and the single-source detectors across both binary and multiclass classification tasks. This highlights the strong generalization performance of VLMs trained on large-scale corpora in Lrq-Fact, positioning it as a potential baseline framework for future multimodal misinformation detection research.
Table 2: F1-score of single-source detectors on both binary and multi-class misinformation detection tasks for MMFakeBench. TVD, VVD, and CCD represent textual veracity distortion, visual veracity distortion, and cross-modal consistency distortion, respectively. The last highlighted row represents the result of Lrq-Fact.
Model Name Train Source Binary F1$\uparrow$ Multiclass F1$\uparrow$
FakingFakeNews [21] TVD 37.8 -
CNNSpot [76] VVD 23.8 -
UnivFD [52] VVD 28.9 -
LNP [40] VVD 33.0 -
FakeNewsGPT4 [44] CCD 41.7 -
HAMMER [61] CCD 43.0 -
LLaVA-1.6-34B [42] MMFakeBench (N=1000) 62.9 25.7
GPT-4V-1.7T [53] MMFakeBench (N=1000) 72.3 51.0
Paligemma-3B [6] MMFakeBench (N=1000) 76.2 63.3
Figure 3: A case study in multimodal misinformation with cross-modal distortion.
4.5.3 Comparison of Different Categories.
We compare the F1 scores of different VLMs across various misinformation types, including Real, Textual Veracity Distortion (TVD), Visual Veracity Distortion (VVD), and Cross-modal Consistency Distortion (CCD), as shown in Table 3. Across all categories, most open-source models significantly underperform compared to Lrq-Fact and GPT-4V, with the largest gap observed in the handling of textual veracity distortion. This highlights the particular difficulty that open-source models face in accurately detecting textual misinformation.
Table 3: F1-score of different VLMs across various misinformation categories: Real, TVD, VVD, and CCD on MMFakeBench. The last highlighted row represents the result of Lrq-Fact.
@@ 4.5 Experiment Results
VILA-13B [39] 32.4 13.4 4.3 37.6 21.9
InstructBLIP-13B [11] 41.9 18.8 19.6 23.8 26.0
BLIP2-FLAN-T5-XXL [34] 41.5 39.2 13.1 32.6 31.6
LLaVA-1.6-34B [42] 51.1 37.6 61.7 49.2 49.9
GPT-4V-1.7T [53] 65.3 67.2 57.3 56.5 61.6
Paligemma-3B [6] 66.8 69.2 56.1 66.2 63.3
Case Study. To demonstrate the effectiveness of our method more directly, we showcase an example where our method successfully detects and classifies the multimodal misinformation in Figure 3. More examples can be found in Appendix (see Figure 4, 5, and 6).
5 Conclusion and Future Work
This paper introduces a novel framework, Lrq-Fact, which effectively leverages LLMs and VLMs to enhance multimodal fact-checking. By generating comprehensive, multi-aspect questions for both image and text content, the framework addresses the challenges of detecting misinformation across different modalities. Extensive experiments conducted on two benchmark datasets, MMFakeBench and DGM4, demonstrate that Lrq-Fact significantly outperforms existing methods in terms of detection accuracy and rationale quality. Through its rule-based decision-making module, the framework automates the fact-checking process and provides clear, transparent justifications for its final decision.
Despite these promising results, several directions for future work remain open. First, incorporating external resources, such as Google Search, to retrieve up-to-date information could further enhance the fact-checking process [2, 65]. Second, expanding the analysis of images to include more thorough investigations, like verifying the authenticity of famous individuals depicted, would improve detection capabilities. Third, exploring alternative approaches to ensure the factuality of generated answers, including checking the answers themselves [74, 17], could strengthen the framework’s reliability. Moreover, explorations on the robustness [30, 77, 78, 67, 13] of the system is also important.
A Appendix
A.1 Evaluation Results for Binary Classification.
In addition to multi-class classification, we also assess the binary classification performance to evaluate the overall detection capability of the baseline models across the datasets. Based on the four categories used in the multimodal misinformation detection task, we map these into binary labels. Specifically, instances labeled as “textual veracity distortion”, “visual veracity distortion”, and “crossmodal consistency distortion” are grouped under the "Fake" category, while the label "True" is used to represent real data. Similar to the multi-class evaluation, we employ the widely used F1 score as the primary metric. In addition, precision, recall, and accuracy are reported as supplementary metrics.
Table 4: Performance comparison (%) of different models for binary classification on the MMFakeBench and DGM4 datasets. The last two highlighted rows represent the results from our proposed framework, Lrq-Fact.
Model Language MMFakeBench (1000) DGM4 (1000)

### v2

@@ 5.4 Rule-Based Decision-Maker
To enhance the judge LLM’s ability to effectively utilize multimodal evidence and make accurate predictions, we introduce a rule-based decision-maker module. This module guides the judge LLM to follow expert-like reasoning steps:
  • General Instructions: We provide guidelines to help judge LLMs establish connections between QA analyses and specific labels. For example, if evidence from FCQs contradicts a claim, the decision-maker assigns it the label “TVD”.
  • Additional Guidelines: We outline supplementary instructions, such as analyzing facial expressions, identifying unrealistic elements, and verifying cross-modal consistency, enabling judge LLMs to conduct a human-like examination.
  • Output Format: We specify the required output format, including the judgment, a fine-grained misinformation label, and a detailed explanation.
@@
For each piece of analyzed content, the decision-maker provides: (1) a final judgment $j_{i}\in${Real, Fake}, (2) a label $l_{i}\in$Textual Veracity Distortion, Visual Veracity Distortion, Mismatch} identifying the specific issue, and (3) a detailed explanation of the decision-making process $e_{i}$. This rule-based approach ensures that the framework provides clear, evidence-based conclusions, carefully weighing the alignment between textual and visual information to detect misinformation. This process can be formulated as:
$qa_{i}^{v}=\oplus_{j=1}^{m}[ques_{i}^{v_{j}},ans_{i}^{v_{j}}],$ (6)
$qa_{i}^{t}=\oplus_{j=1}^{n}[ques_{i}^{t_{j}},ans_{i}^{t_{j}}],$ (7)
$j_{i},l_{i},e_{i}={\rm LLM}(text_{i},img_{i},des_{i},qa_{i}^{v},qa_{i}^{t}),$ (8)
where $\oplus$ denotes the concatenation operation, and $m$ and $n$ represent the number of visual and textual FCQs, respectively.
@@ 5.5 Datasets
MMFakeBench Liu et al. (2024b) contains 11,000 image-text pairs. It goes beyond the assumption of single-source forgery and presents samples with “Real”, “Textual Veracity Distortion”, “Visual Veracity Distortion”, and “Cross-modal Consistency Distortion”, with both human- and machine-generated images.
DGM4 Shao et al. (2023) is a large-scale multimodal dataset comprising 230,000 image-text paired samples. Image manipulation in the dataset involves “face swapping and facial emotion editing”, while text manipulation includes “sentence replacement and textual sentiment editing”. The DGM4 dataset is constructed based on the VisualNews dataset Liu et al. (2020), which collects data from multiple news agencies.
Factify Mishra et al. (2022) is a multimodal fact-checking benchmark comprising 50,000 data points, each consisting of a textual claim, an associated image, and corresponding reference documents. It is categorized into three main classes: “Support”, “NotEnoughInfo”, and “Refute”, with finer-grained labels for detailed evaluation.
More details about the datasets are provided in Appendix A.5.
Backbone Approach MMFakeBench DGM4 Factify
@@
VILA Lin et al. (2024) SP 11.5 30.0 19.4 19.8 25.6 27.7
InstructBLIP Dai et al. (2023) SP 13.7 28.8 19.2 19.8 24.3 26.9
BLIP-2 Li et al. (2023) SP 16.7 32.8 18.1 19.1 26.6 28.6
LLaVA-1.6 Liu et al. (2024a) SP 25.7 40.4 32.5 39.4 51.3 56.2
GPT-4V-1.7T OpenAI (2023) SP 51.0 54.0 42.3 51.5 64.2 68.2
GPT-4o OpenAI (2023) SP 49.2 60.9 39.9 55.9 72.5 71.2
Lrq-Fact (w/o RAG) FCQs 66.5 65.5 45.8 58.0 - -
Lrq-Fact (w/ RAG) FCQs 71.6 70.8 49.2 62.3 75.2 73.1
Table 2: The last two rows represent the results of Lrq-Fact. Standard prompt (SP) refers to a generic prompt without fact-checking questions. Bold: best result; Underline: second best result.
@@ 5.6 Experiment Details and Settings
For MMFakeBench and DGM4, we sample 1,000 validation instances, ensuring a balanced distribution: 300 Real, 300 TVD, 100 VVD, and 300 CMM. DGM4 labels are mapped as follows: real (“orig”), VVD (“face swap,” “face attribute”), TVD (“text attribute”), and mismatch (“text swap”) to ensure consistency across datasets. For Factify, we sample 750 validation instances: 300 Support (Support Multimodal, Support Text), 300 NotEnoughInfo ( Insufficient Text, Insufficient Multimodal), and 150 Refute, maintaining a balanced evaluation.
@@
Lrq-Fact integrates both LLMs and VLMs to handle the tasks of question generation and answering. For generating visual and textual questions, we use the GPT-4o Achiam et al. (2023), along with Paligemma Beyer et al. (2024) and LLaMA 3.1 Vavekanand and Sam (2024) as different backbones for VLMs and LLMs respectively. For each modality, we generate five questions by employing in-context learning, where 10 example questions are provided for each modality to guide the model in generating high-quality and relevant questions. To answer fact-checking questions, we employ two settings: (1) using only LLM knowledge and (2) a RAG approach, where an external knowledge retrieval module fetches relevant supporting documents before generating responses. In the decision-making phase, GPT-4o is also used to leverage the FCQs for fact verification. Known for its strong reasoning and rule-following capabilities, GPT-4o assesses the veracity of the content and offers clear rationales for its conclusions. All experiments are conducted on NVIDIA 40GB V100 GPUs.
5.6.1 Evaluation Metrics
@@
To assess the performance of the various baselines, we employ a multi-class classification approach. Following the practices established in prior works Zhang and Gao (2023); Qian et al. (2021); Chen et al. (2023), we use the widely adopted macro-F1 score as our primary evaluation metric. The macro-F1 score provides a balanced measure of precision and recall through their harmonic mean, ensuring fair evaluation across all classes. In addition to the macro-F1 score, we also report macro-accuracy as complementary metrics, offering a more comprehensive understanding of model performance.
5.6.2 Comparison Models
@@
We select a diverse range of VLMs as baseline models for comparison. These include: InstructBLIP Dai et al. (2023), VILA Lin et al. (2024), BLIP-2 Li et al. (2023), LLaVA-1.6 Liu et al. (2024a) and the closed-source GPT-4V-1.7T, GPT-4o OpenAI (2023) models.
5.7 Experiment Analysis
To assess whether LLM-generated FCQs enhance multimodal fact-checking, we conduct a three-stage analysis: (1) a comparison between our Lrq-Fact and other baseline methods, (2) an ablation study to measure the impact of generated questions on fact-checking performance and (3) an in-depth examination of how different types of questions contribute to improving verification accuracy across modalities.
Method MMFakeBench DGM4 Factify
F1 ACC F1 ACC F1 ACC
Lrq-Fact (VLM: Paligemma & LLM: LLaMA 3.1) 51.2 56.4 41.2 53.6 66.2 65.3
     w/ Textual FCQs 62.5+22.1% 59.1+4.8% 43.3+5.1% 54.1+0.9% 64.8-2.1% 60.9-6.7%